# Scalable Private Search with Wally

### v4

@@ 2 Background
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [36, Lemma 2.3].
@@
Though the paper is mostly presented for approximate nearest neighbor retrieval, Wally applies to exact retrieval as well by using keyword PIR for its homomorphic computation. Private information retrieval (PIR) protocols are cryptographic client-server protocols where a server holds a large database and the client obliviously retrieves an entry from the database. PIR protocols can be for index retrieval or retrieval by keyword [12]. In this paper, we focus on single-server PIR protocols [30] since we can shard the database and the client can run Wally to anonymously retrieve entries and hide its downloaded entry on the shard. We use and optimize PIR protocols based on SHE with minimal preprocessing, such as [5] and [4]. There is a notion of differentially-private PIR in the multi-server setting [34, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy. This is in contrast to Wally which stores the database on one server or servers controlled by a single entity.
@@ Appendix A Optimizations
Here we present our optimizations for both use cases: approximate nearest neighbor search and exact retrieval (PIR). Our optimizations revolve around the BFV [9, 18] and BGV [10] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [4]. First, we describe the standard RNS optimization in SHE [21] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [28].
@@
We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$. Concretely, we are able to drop around $8$kB from the response size in approximate nearest neighbors search to a response size of $47$kB when combined with our BFV plaintext CRT optimization described in the next subsection.
A.1 Approximate Nearest Neighbors Optimizations
The homomorphic computation in the approximate nearest neighbors use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size. We scale $\mathbf{e}$’s entries to integers, round them to a vector $\mathbf{\widetilde{e}}$, and perform the inner product over the integers. The computation is correct as long as the inner product does not wrap around mod $t$. We define plaintext precision as $\approx\log_{2}(t)/2$ since the inner product has a multiplicative depth of one.
@@ A.2 PIR Optimizations
We apply the lazy rescaling technique of Kim et al [29] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [29, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.

### v5

@@ 2 Background
Essentially, standard DP ensures that neighboring datasets have a privacy loss of at most $\epsilon$ with probability at least $1-\delta$ for all observations. We use basic composition in DP [35, Lemma 2.3].
@@
Though the paper is mostly presented for semantic search, Wally applies to keyword search as well by using keyword variant of private information retrieval (PIR). PIR protocols are cryptographic client-server protocols where a server holds a large database and the client can privately retrieves an entry corresponding to a keyword from the database. In this paper, we focus on protocols based on a single server [29]. There is a also notion of differentially-private PIR in the multi-server setting [33, 2], where the database is stored across non-colluding servers while guaranteeing differential privacy. We can not directly use these protocols in Wally which stores the database on one server or multiple servers controlled by a single entity.
@@ Appendix A Optimizations
Here we present our optimizations for both use cases: private nearest neighbor search and exact retrieval (PIR). Our optimizations revolve around the BFV [9, 17] and BGV [10] SHE schemes. Further, we describe several optimizations to the MulPIR scheme [4]. First, we describe the standard RNS optimization in SHE [20] and a method to compress ciphertexts by dropping LSBs as described in Cheetah [27].
@@
We have correctness as long as $\|e+e_{drop}\|_{\infty}<Q^{\prime}/2t$. Concretely, we are able to drop around $8$kB from the response size in private nearest neighbor search to a response size of $47$kB when combined with our BFV plaintext CRT optimization described in the next subsection.
A.1 Private Nearest Neighbor Search Optimizations
The homomorphic computation in the private nearest neighbor search use case is $C$ inner products over real embedding vectors satisfying $\mathbf{e}$, normalized to $\|\mathbf{e}\|_{2}=1$ in our use case, where $C$ is the maximum cluster size. We scale $\mathbf{e}$’s entries to integers, round them to a vector $\mathbf{\widetilde{e}}$, and perform the inner product over the integers. The computation is correct as long as the inner product does not wrap around mod $t$. We define plaintext precision as $\approx\log_{2}(t)/2$ since the inner product has a multiplicative depth of one.
@@ A.2 PIR Optimizations
We apply the lazy rescaling technique of Kim et al [28] in MulPIR’s last step, the inner product between ciphertexts $\mathbf{r}^{T}\mathbf{a}$. The main idea here is that BFV multiplications first multiply polynomials over the integers, then scale and round back to integers modulo $Q$. Kim et al [28, Appendix F] noticed that one can multiply over the integers, add over the integers, then scale and round after the additions, from the dimension number of scaling operations to one. Lazy rescaling yielded a $15-20\%$ improvement in our MulPIR implementation.